15x+11y= 132
2x+3y=59

Answer by Alan3354(69443)

15x+11y= 132
2x+3y=59
--------------
|15 11 -132|
|+2 +3 -59|
-----------
Det = 15*3 - 2*11 = 23
------
Dx = 11*(-59) - 3*(-132) = -253
x = -253/23 = -11
------
-Dy = 15*(-59) - 2*(-132) = -621
y = 621/23 = 27